# Cryptanalysis of PLWE based on zero-trace quadratic roots

### v1

@@ 3.1. Attack developed by [36] and [37]
Proposition 3.1 ([36], Proposition 3.1).
Assume $|\Sigma|<q$ and let $M$ be the number of input samples. If Algorithm 1 returns NOT PLWE, then the samples come from the uniform distribution. If it outputs anything other than NOT PLWE, then the samples are valid PLWE samples with probability $1-\left(|\Sigma|/q\right)^{M}$. In particular, this probability tends to 1 as $M$ grows.
@@
Notice that the cyclotomic polynomial $\Phi_{n}(x)$ is protected against these attacks. Indeed, $\alpha=1$ is never a root modulo $q\nmid n$. Moreover, for $q\equiv 1\pmod{n}$ the order of each of the $\phi(n)$ different roots of $\Phi_{n}(x)$ is precisely $n$.
@@
In this subsection, we present an attack against PLWE for a ring $R_{q}:=\mathbb{F}_{q}[x]/(f(x))$ with $q$ prime and $f(x)=\sum_{j=0}^{N}f_{j}x^{j}\in\mathbb{Z}[x]$ monic and irreducible over $\mathbb{Z}[x]$ of degree $N$. We start our study by assuming, first, that $x^{2}+1$ is irreducible over $\mathbb{F}_{q}[x]$, and we assume too that $f(x)$ is divisible by $x^{2}+1$ in $\mathbb{F}_{q}[x]$. Second, we will study the case where $f(x)$ has an irreducible quadratic factor of the form $x^{2}+\rho$ with $\rho\in\mathbb{F}_{q}[x]$ with suitable multiplicative order modulo $q$.
@@
Suppose first that $\left(\frac{-1}{q}\right)=-1$ and that $x^{2}+1\mid f(x)$ in $\mathbb{F}_{q}[x]$. Hence, there exists $\alpha\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ with $\alpha^{2}=-1$ and $f(\alpha)\equiv 0\pmod{q}$. Notice that $Tr(\alpha)=0$, where $Tr$ stands for the trace of $\mathbb{F}_{q^{2}}$ over $\mathbb{F}_{q}$.
@@
Now, if $(a(x),b(x)=a(x)s(x)+e(x))\in R_{q}^{2}$ is a PLWE sample attached to a secret $s(x)$ and an error term $e(x)=\sum_{i=0}^{N-1}e_{i}x^{i}$, where the coefficients $e_{i}$ are sampled from independent discrete Gaussians of mean 00 and variance $\sigma^{2}$, then
$b(\alpha)-a(\alpha)s=e(\alpha),$
with $s:=s(\alpha)\in\mathbb{F}_{q^{2}}$ and
(3.1) $Tr(b(\alpha)-a(\alpha)s)=Tr(e(\alpha))=\sum_{j=0}^{N-1}e_{j}t_{j},$
@@
where $t_{j}=Tr(\alpha^{j})$. But if $j=2j^{\prime}$, then $t_{j}=Tr((-1)^{j^{\prime}})=2(-1)^{j^{\prime}}$, otherwise $t_{j}=0$. Hence we are left with:
(3.2) $Tr(b(\alpha)-a(\alpha)s)=2\sum_{j=0}^{N^{*}}e_{2j}(-1)^{j},$
where $N^{*}:=\left\lfloor\frac{N-1}{2}\right\rfloor$. The sum of errors in Equation (3.2) is sampled from a discrete Gaussian $\mathcal{N}\left(0,\sqrt{N^{*}}\sigma\right)$, and we can list those elements which occur with probability beyond $0.95$, namely, the integer values in the interval
(3.3) $\Sigma:=[-2\sqrt{N^{*}}\sigma,2\sqrt{N^{*}}\sigma],$
@@
containing at most $4\sqrt{N^{*}}\sigma+1=\mathcal{O}(q^{1/4})$ elements, which we will term the smallness region for this attack.
@@
More in general, suppose now that $\left(\frac{-\rho}{q}\right)=-1$ and denote by $r$ the multiplicative order of $-\rho$ modulo $q$. As before, assume that $x^{2}+\rho\mid f(x)$ and that $f(\alpha)=0$ with $\alpha\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ and $\alpha^{2}=-\rho$. In this case, for a PLWE-sample $(a(x),b(x)=a(x)s(x)+e(x))$ with $e(x)=\sum_{i=0}^{N-1}e_{i}x^{i}$, we have for $s=s(\alpha)$:
(3.4) $Tr(b(\alpha)-a(\alpha)s)=Tr(e(\alpha))=2\sum_{k=0}^{r-1}(-\rho)^{k}\sum_{i=0}^{\left\lfloor\frac{N^{*}}{r}\right\rfloor-1}e_{2(ir+k)},$
In this case, for each $10\leq k\leq r-$ 1 the sum of coefficients $\sum_{i=0}^{\left\lfloor\frac{N^{*}}{r}\right\rfloor-1}e_{2(ir+k)}$ are sampled from a centred discrete Gaussian of standard deviation less than or equal to $\sqrt{\left\lfloor\frac{N^{*}}{r}\right\rfloor}\sigma$, hence we can list those elements which occur with probability beyond $0.95$, which are at most $14\sqrt{\left\lfloor\frac{N^{*}}{r}\right\rfloor}\sigma+$ 1. Next, we can construct a look-up table $\Sigma$ for all the possible values in Equation (3.4), which occur with probability beyond $(0.95)^{r}$. Notice that
(3.5) $|\Sigma|\leq\left(4\sqrt{\left\lfloor\frac{N^{*}}{r}\right\rfloor}\sigma+1\right)^{r}.$
@@
What makes suitable the order of $-\rho$ to derive an attack is not (just) being small, but making the upper bound of Equation (3.5) much smaller than $q$, as we will analyse in the rest of the Section. We turn the previous ideas into an attack but first, for a reason which will be transparent later, we define the following set:
@@
The set $R_{q,0}$ is a ring, which, in particular, as an $\mathbb{F}_{q}$-vector subspace of $R_{q}$ has dimension $N-1$.
@@
Let us set $S:=\mathbb{F}_{q^{2}}$, and assume that we are given a set of samples from $R_{q,0}\times R_{q}$. The goal is to distinguish whether these samples come from the PLWE distribution (but with the $a(x)$ term belonging to $R_{q,0}$) or from a uniform distribution with values in $R_{q,0}\times R_{q}$. In the sequel, we provide an algorithm that achieves that goal for the particular case in which $\rho=1$, though it is immediate to convert it to the general case.
In summary, the algorithm runs as follows: given a sample $(a_{i}(x),b_{i}(x))\in R_{q,0}\times R_{q}$, we pick a guess $s\in S$ for $s(\alpha)$ and check whether $\frac{1}{2}Tr(b_{i}(\alpha)-a_{i}(\alpha)s)$ belongs to $\Sigma$. If this is not the case, we can safely remove from $S$ not only $s$, but also all the elements $s^{\prime}\in\mathbb{F}_{q^{2}}$ such that $Tr(s^{\prime})=Tr(s)$. This means that we can remove $q$ elements of $S$ at a time.
Since $a(\alpha)\in\mathbb{F}_{q}$, then
$\frac{1}{2}Tr(b(\alpha)-a(\alpha)s)=\frac{1}{2}Tr(b(\alpha))-\frac{1}{2}a(\alpha)Tr(s).$
Therefore, it is enough to check, for each $g\in\mathbb{F}_{q}$ (so that $g$ is a putative value for $Tr(s)$), whether or not
$\frac{1}{2}Tr(b(\alpha))-\frac{1}{2}a(\alpha)g\in\Sigma.$
This is at the price that if the algorithm returns an element $g\in\mathbb{F}_{q}$, we should understand that this element is just the trace of one of the $q$ possible guesses for $s(\alpha)$. However, this is (even if weaker than Algorithm 1) enough as a decision attack.
@@
As pointed out above, notice that if $|G|=1$, say $G=\{g\}$, unlike Algorithm 1, our Algorithm 2 does not output a guess for $s(\alpha)$; instead, $g$ is a guess for $\frac{1}{2}Tr(s(\alpha))$ and in this case, we can venture that there exists $s(x)\in R_{q}$ such that $b_{i}(x)-a_{i}(x)s(x)$ belongs to the error distribution for every sample $(a_{i}(x),b_{i}(x))$.
@@
  • For each sample $(a_{i}(x),b_{i}(x))$, evaluating $Tr(b_{i}(\alpha))$, requires no multiplications in $\mathbb{F}_{q}$. Hence checking whether $(Tr(b(\alpha))-a(\alpha)g)/2\in\Sigma$ requires only $1$ multiplication in $\mathbb{F}_{q}$.
@@
$(Tr(b(\alpha))-a(\alpha)g)/2\in\Sigma\mbox{ if and only if }b^{*}(-\rho)-ga^{*}(-\rho)\in\Sigma.$
@@
Assume that $|\Sigma|<q$ and let $M$ be the number of input samples. If Algorithm 2 returns NOT PLWE, then the samples come from the uniform distribution on $R_{q,0}\times R_{q}$. If it outputs anything other than NOT PLWE, then the samples are valid PLWE samples with probability $1-\left(|\Sigma|/q\right)^{M}$. In particular, this probability tends to 1 as $M$ grows.
@@
Let us denote by $X$ an unknown random distribution. The target of Algorithm 2 is to determine whether $X$ is a uniform or a PLWE distribution. For a sample $v=(a(x),b(x))\in R_{q,0}\times R_{q}$ output by $X$, we can write
@@
Notice that if $X$ is a uniform distribution, then Lemma 3.6 ensures that $X^{*}$ is also a uniform distribution with values on $R_{q}^{*}\times R_{q}^{*}$. As for the reciprocal, suppose that $X^{*}$ is uniform, but $X$ is PLWE with a Gaussian distribution $\mathcal{N}(0,\sigma^{\ast})$ associated to the error. Then consider pairs $(a_{i}^{\ast}(x),b_{i}^{\ast}(x))$ output by $X^{\ast}$; since by hypothesis $X^{\ast}$ is uniform, evaluating $(a_{i}^{*}(x),b_{i}^{*}(x))$ at $-\rho$ should yield a uniform distribution of values, by virtue of Lemma 3.6 and Lemma 2.2; however, that evaluation equals $(a_{i}^{*}(-\rho),a_{i}^{*}(-\rho)s^{*}(-\rho)+e_{i}^{*}(-\rho))$, where each $a_{i}^{*}(-\rho)$ is indeed uniformly sampled due to Proposition 2.3 and Lemma 2.2; but each $e_{i}^{*}(-\rho)$ is the output of the Gaussian associated to $X$, which we assumed PLWE. Thus, $X^{\ast}$ turns out to be not uniform, against the hypothesis. In summary, the distribution $X$ is uniform if and only if $X^{*}$ is uniform.
If $X$ is the PLWE distribution, for a sample $(a(x),b(x))$ it will be $b(x)=a(x)s(x)+e(x)$ with $e(x)$ drawn from an $R_{q}$-valued discretised Gaussian, hence setting $a(x)=a_{0}(x)+a_{1}(x)$, and likewise for $b(x)$, $s(x)$ and $e(x)$ we have that the even degree part of $b(x)$ is
@@
Assume now that $X$ is a PLWE distribution. Any $v=(a(x),b(x))$ sampled from $X$ has $a(x)$ sampled from an $R_{q}$-valued uniform distribution and $b(x)=a(x)s(x)+e(x)$, where $e(x)$ is sampled from $\chi_{R}$, an $R_{q}$-valued discrete Gaussian variable. Taking into account that for any $i\geq 1$, it holds that
$P[C_{i,v}]=P[X=v]=P[U=a(x)\cap\chi_{R}=e(x)],$
@@
Within $X_{0}$, the variable $count$ stores the expected number of times one has to run $X$ before succeeding. Since this random variable is distributed as a geometric random variable of first kind with success probability $q^{-1}$, the expected number of times is just its mean $E[count]=q=\mathcal{O}(N^{2})$ (see for instance [12, Appendix A]).
@@
Next, we must choose $k$ so that $1-\left(|\Sigma|/q\right)^{k}\geq\theta$, where $\theta$ is our desired success probability threshold.

### v3

@@ 3.1. Attack developed by [13] and [14]
Proposition 3.1.
Let $\Sigma$, $M$ and $r$ be defined as above, and assume $|\Sigma|<q$. Then
  1. (1)
    If Algorithm 1 returns PLWE or NOT ENOUGH SAMPLES, then the samples are PLWE with probability at least $1-q\left(\frac{|\Sigma|}{q}\right)^{M}$.
  2. (2)
    If Algorithm 1 returns NOT PLWE, then the samples are uniform with probability $1$.
Proof.
  1. (1)
    Let $E$ represent the event “Algorithm 1 returns PLWE or NOT ENOUGH SAMPLES”. By using Bayes’ theorem we have
    $\mathrm{P}(D=\mathcal{G}_{\sigma}|E)=1-\mathrm{P}(D=U|E)$
    $=1-\frac{\mathrm{P}(E|D=U)\cdot\mathrm{P}(D=U)}{\mathrm{P}(E|D=U)\cdot\mathrm{P}(D=U)+\mathrm{P}(E|D=\mathcal{G}_{\sigma})\cdot\mathrm{P}(D=\mathcal{G}_{\sigma})}$
    (3.1) $=1-\frac{\mathrm{P}(E|D=U)}{\mathrm{P}(E|D=U)+\mathrm{P}(E|D=\mathcal{G}_{\sigma})}.$
    Notice that we are implicitly assuming that both distributions can be chosen with equal probability.
    Observe that $\mathrm{P}(E|D=\mathcal{G}_{\sigma})$ is the probability that there exist some $g\in\mathbb{F}_{q}$ such that $b_{i}(\alpha)-a_{i}(\alpha)g\in\Sigma$, for all $i=1,\dotsc,M$. Since we are supposing that the distribution is precisely $\mathcal{G}_{\sigma}$, then at least one $g$ is sure to exist, namely, $s(\alpha)$, the evaluation of the secret. Then we have that $\mathrm{P}(E|D=\mathcal{G}_{\sigma})\geq\mathrm{P}(s(\alpha)\in\Sigma)=1$.
    To compute $\mathrm{P}(E|D=U)$, denote by $U_{g}$ the event that $b_{i}(\alpha)-a_{i}(\alpha)g\in\Sigma$, for all $i=1,\dotsc,M$. Then
    $\mathrm{P}(E|D=U)=\mathrm{P}\left(\bigcup_{g=0}^{q-1}U_{g}\right)\leq\sum_{g=0}^{q-1}\mathrm{P}(U_{g}),$
    as the events $U_{g}$ might not be independent. But we are now supposing that the values $a_{i}$ and $b_{i}$ are uniform, thus $b_{i}-a_{i}g$ is also uniform due to Lemma 2.2. Hence we have that $\mathrm{P}(b_{i}-a_{i}g\in\Sigma)=|\Sigma|/q$ for any $i$ and any $g$, so
    $\mathrm{P}(U_{g})=\prod_{i=1}^{M}\frac{|\Sigma|}{q}=\left(\frac{|\Sigma|}{q}\right)^{M},$
    and henceforth one obtains:
    $\mathrm{P}(E|D=U)\leq q\left(\frac{|\Sigma|}{q}\right)^{M}.$
    Setting all these values into Equation (1), we eventually arrive at:
    $\mathrm{P}(D=\mathcal{G}_{\sigma}|E)=1-\frac{\mathrm{P}(E|D=U)}{\mathrm{P}(E|D=U)+1}\geq 1-\mathrm{P}(E|D=U)\geq 1-q\left(\frac{|\Sigma|}{q}\right)^{M}.$
  2. (2)
    Let $E^{\prime}$ stand for event “Algorithm 1 returns NOT PLWE”. Again using Bayes’ theorem we can write:
    $\mathrm{P}(D=U|E^{\prime})$ $=$ $\frac{\mathrm{P}(E^{\prime}|D=U)\cdot\mathrm{P}(D=U)}{\mathrm{P}(E^{\prime}|D=U)\cdot\mathrm{P}(D=U)+\mathrm{P}(E^{\prime}|D=\mathcal{G}_{\sigma})\cdot\mathrm{P}(D=\mathcal{G}_{\sigma})}$
    $=$ $\frac{\mathrm{P}(E^{\prime}|D=U)}{\mathrm{P}(E^{\prime}|D=U)+\mathrm{P}(E^{\prime}|D=\mathcal{G}_{\sigma})}.$
    But $\mathrm{P}(E^{\prime}|D=\mathcal{G}_{\sigma})=1-\mathrm{P}(E|D=\mathcal{G}_{\sigma})=0$, hence $\mathrm{P}(D=U|E^{\prime})$, as required.
@@
We point out that the conclusion of [13, Proposition 3.1] is slightly different to ours. Namely, our conclusion when Algorithm 1 returns NOT PLWE is exactly the same, but in the case when Algorithm 1 returns PLWE or NOT ENOUGH SAMPLES our upper bound is slightly weaker. However, we think that in the proof provided in [13, Proposition 3.1] the probability $P(E|D=U)$ is underestimated. Moreover, in [13] it is implicitly assumed that the sum of $N$ independent discrete truncated Gaussian distributions of variance $\sigma^{2}$ is a discrete truncated Gaussian distribution of variance $N\sigma^{2}$, as in the continuous non-truncated case. This claim needs further clarification not provided therein (see, for instance [15]). We have avoided this approach just considering that our smallness regions consist of all the possible values for the error polynomials evaluated at the roots.
Remark 3.3.
Notice that these attacks do not apply to cyclotomic moduli $\Phi_{n}(x)$. Indeed, $\alpha=1$ is never a root modulo $q\nmid n$. Moreover, for $q\equiv 1\pmod{n}$ the order of each of the $\phi(n)$ different roots of $\Phi_{n}(x)$ is precisely $n$.
@@
In this subsection, we present an attack against PLWE for a ring $R_{q}:=\mathbb{F}_{q}[x]/(f(x))$ with $q$ prime and $f(x)=\sum_{j=0}^{N}f_{j}x^{j}\in\mathbb{Z}[x]$ monic and irreducible over $\mathbb{Z}[x]$ of degree $N$. For the rest of this section, let us define $N^{*}:=\left\lfloor\frac{N-1}{2}\right\rfloor$.
@@
We start our study by assuming, first, that $x^{2}+1$ is irreducible over $\mathbb{F}_{q}[x]$, and that $f(x)$ is divisible by $x^{2}+1$ in $\mathbb{F}_{q}[x]$. Second, we will study the case where $f(x)$ has an irreducible quadratic factor of the form $x^{2}+\rho$ with $\rho\in\mathbb{F}_{q}[x]$ with suitable multiplicative order modulo $q$.
@@
Suppose first that $x^{2}+1$ is irreducible in $\mathbb{F}_{q}[x]$ and that $x^{2}+1\mid f(x)$ in $\mathbb{F}_{q}[x]$. Hence, there exists $\alpha\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ with $\alpha^{2}=-1$ and $f(\alpha)\equiv 0\pmod{q}$. Notice that $\operatorname{Tr}(\alpha)=0$, where $\operatorname{Tr}$ stands for the trace of $\mathbb{F}_{q^{2}}$ over $\mathbb{F}_{q}$.
@@
Now, if $(a(x),b(x)=a(x)s(x)+e(x))\in R_{q}^{2}$ is a PLWE sample attached to a secret $s(x)$ and an error term $e(x)=\sum_{i=0}^{N-1}e_{i}x^{i}$, where the coefficients $e_{i}$ are sampled from independent discrete Gaussians of mean 0 and variance $\sigma^{2}$, then
$b(\alpha)-a(\alpha)s(\alpha)=e(\alpha),$
where now $s(\alpha)\in\mathbb{F}_{q^{2}}$, and
(3.2) $\operatorname{Tr}(b(\alpha)-a(\alpha)s(\alpha))=\operatorname{Tr}(e(\alpha))=\sum_{j=0}^{N-1}e_{j}t_{j},$
@@
where $t_{j}=\operatorname{Tr}(\alpha^{j})$. But if $j=2j^{\prime}$, then $t_{j}=\operatorname{Tr}((-1)^{j^{\prime}})=2(-1)^{j^{\prime}}$, otherwise $t_{j}=0$. Hence we are left with:
(3.3) $\operatorname{Tr}(b(\alpha)-a(\alpha)s(\alpha))=2\sum_{j=0}^{N^{*}}e_{2j}(-1)^{j}.$
The sum of errors in Equation (3.3) takes values on the set $4N^{*}\sigma,4N^{*}\sigma]\cap\mathbb{Z}$. Hence we can list the possible values $\frac{1}{2}\left(\operatorname{Tr}(b(\alpha)-a(\alpha)s(\alpha))\right)$. These elements are contained in the set
(3.4) $\Sigma:=[-2N^{*}\sigma,2N^{*}\sigma]\cap\mathbb{Z},$
@@
containing at most $4N^{*}\sigma+1$ elements, which we will term the smallness region for this attack.
@@
More in general, suppose now that for some non-zero $\rho\in\mathbb{F}_{q}$, the polynomial $x^{2}+\rho$ is irreducible over $\mathbb{F}_{q}[x]$ and denote by $r$ the multiplicative order of $\rho$ modulo $q$. As before, assume that $x^{2}+\rho\mid f(x)$ and that $\alpha^{2}=-\rho$ with $\alpha\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ so that $f(\alpha)\equiv 0\pmod{q}$. In this case, for a PLWE-sample $(a(x),b(x)=a(x)s(x)+e(x))$ with $e(x)=\sum_{i=0}^{N-1}e_{i}x^{i}$, we have:
(3.5) $\operatorname{Tr}(b(\alpha)-a(\alpha)s(\alpha))=\operatorname{Tr}(e(\alpha))=2\sum_{k=0}^{r-1}(-\rho)^{k}\epsilon_{k}$
with $\epsilon_{k}:=\sum_{i=0}^{\left\lfloor\frac{N^{*}}{r}\right\rfloor-1}e_{2(ir+k)}$, for $0\leq k\leq r-1$.
Now each coefficient $\epsilon_{k}$ takes values on the set $\left[2\frac{N^{*}}{r}\sigma,2\frac{N^{*}}{r}\sigma\right]\cap\mathbb{Z}$. So, let us denote again by $\Sigma$ the set of all the possible values of $\frac{1}{2}\operatorname{Tr}(e(\alpha))$. We have that
(3.6) $|\Sigma|\leq\left(4\frac{N^{*}}{r}\sigma+1\right)^{r}.$
@@
What makes suitable the order of $\rho$ to derive an attack is not (just) being small, but making the upper bound of Equation (3.6) much smaller than $q$, as we will analyse in the rest of the Section. We turn the previous ideas into an attack but first, for a reason which will be transparent later, we define the following set:
@@
The set $R_{q,0}$ is a ring, which, in particular, as an $\mathbb{F}_{q}$-vector subspace of $R_{q}$ has dimension N1N-1.
@@
Let us set $S:=\mathbb{F}_{q^{2}}$, and assume that we are given a set of samples from $R_{q,0}\times R_{q}$. The goal is to distinguish whether these samples come from the PLWE distribution (but with the $a(x)$ term belonging to $R_{q,0}$) or from a uniform distribution with values in $R_{q,0}\times R_{q}$.
We provide Algorithm 2 to achieve such goal that, in summary, runs as follows: given a sample $(a_{i}(x),b_{i}(x))\in R_{q,0}\times R_{q}$, we pick a guess $s\in S$ for $s(\alpha)$ and check whether $\frac{1}{2}\operatorname{Tr}(b_{i}(\alpha)-a_{i}(\alpha)s)$ belongs to $\Sigma$. Since $a(\alpha)\in\mathbb{F}_{q}$, then
$\frac{1}{2}\operatorname{Tr}(b(\alpha)-a(\alpha)s)=\frac{1}{2}\operatorname{Tr}(b(\alpha))-\frac{1}{2}a(\alpha)\operatorname{Tr}(s).$
Therefore, it is enough to check, for each $g\in\mathbb{F}_{q}$ (so that $g$ is a putative value for $\operatorname{Tr}(s)$), whether or not
$\frac{1}{2}\operatorname{Tr}(b(\alpha))-\frac{1}{2}a(\alpha)g\in\Sigma.$
If this is not the case, we are actually removing from $S$ all the elements $s^{\prime}\in\mathbb{F}_{q^{2}}$ such that $\operatorname{Tr}(s^{\prime})=g$. This means that we remove $q$ elements from $S$ at a time.
Remark that if the algorithm returns an element $g\in\mathbb{F}_{q}$, we should understand that this element is just the trace of one of the $q$ possible guesses for $s(\alpha)$. However, this is (even if weaker than Algorithm 1) enough as a decision attack.
@@
A lookup table $\Sigma$ of all possible values for $\operatorname{Tr}(e(\alpha))$
@@
  • set $G:=\emptyset$
  • for $g\in\mathbb{F}_{q}$ do
    • for $(a_{i}(x),b_{i}(x))\in C$ do
        if $\frac{1}{2}\left(\operatorname{Tr}(b(\alpha))-a(\alpha)g\right)\notin\Sigma$ then
          next $g$
    • set $G:=G\cup\{g\}$
  • if $G=\emptyset$ then return NOT PLWE
  • if $|G|=1$ then return PLWE
  • if $|G|>1$ then return NOT ENOUGH SAMPLES
@@
Actually, if the output of Algorithm 2 is such that $|G|=1$, say $G=\{g\}$, Algorithm 2 (unlike Algorithm 1) does not output a guess for $s(\alpha)$; instead, $g$ is a guess for $\frac{1}{2}\operatorname{Tr}(s(\alpha))$ and in this case, we can venture that there exists $s(x)\in R_{q}$ such that $b_{i}(x)-a_{i}(x)s(x)$ belongs to the error distribution for every sample $(a_{i}(x),b_{i}(x))$.
@@
  • For each sample $(a_{i}(x),b_{i}(x))$, evaluating $\operatorname{Tr}(b_{i}(\alpha))$, requires no multiplications in $\mathbb{F}_{q}$. Hence checking whether $(\operatorname{Tr}(b(\alpha))-a(\alpha)g)/2\in\Sigma$ requires only $1$ multiplication in $\mathbb{F}_{q}$.
@@
$(\operatorname{Tr}(b(\alpha))-a(\alpha)g)/2\in\Sigma\mbox{ if and only if }b^{*}(-\rho)-ga^{*}(-\rho)\in\Sigma.$
@@
Assume that $|\Sigma|<q$ and let $M$ be the number of input samples. If Algorithm 2 returns NOT PLWE, then the samples come from the uniform distribution on $R_{q,0}\times R_{q}$. If it outputs anything other than NOT PLWE, then the samples are valid PLWE samples with probability $1-q\left(|\Sigma|/q\right)^{M}$. In particular, this probability tends to 1 as $M$ grows.
@@
Let us denote by $X$ an unknown random distribution which can be either uniform or a PLWE one. The target of Algorithm 2 is to determine which is the case. For a sample $v=(a(x),b(x))\in R_{q,0}\times R_{q}$ output by $X$, we can write
@@
Notice that if $X$ is a uniform distribution, then Lemma 3.7 ensures that $X^{*}$ is also a uniform distribution with values on $R_{q}^{*}\times R_{q}^{*}$. As for the reciprocal, suppose that $X^{*}$ is uniform, but $X$ is PLWE with a truncated Gaussian distribution associated to the error. Then consider pairs $(a_{i}^{\ast}(x),b_{i}^{\ast}(x))$ output by $X^{\ast}$; since by hypothesis $X^{\ast}$ is uniform, evaluating $(a_{i}^{*}(x),b_{i}^{*}(x))$ at $\rho$ should yield a uniform distribution of values, by virtue of Lemma 3.7 and Lemma 2.2; however, that evaluation equals $(a_{i}^{*}(-\rho),a_{i}^{*}(-\rho)s^{*}(-\rho)+e_{i}^{*}(-\rho))$, where each $a_{i}^{*}(-\rho)$ is indeed uniformly sampled due to Proposition 2.3 and Lemma 2.2; but each $e_{i}^{*}(-\rho)$ is the output of the truncated Gaussian associated to $X$, which we assumed PLWE. Thus, $X^{\ast}$ turns out to be not uniform, against the hypothesis. In summary, the distribution $X$ is uniform if and only if $X^{*}$ is uniform.
If $X$ is the PLWE distribution, for a sample $(a(x),b(x))$ it will be $b(x)=a(x)s(x)+e(x)$ with $e(x)$ drawn from an $R_{q}$-valued discrete truncated Gaussian, hence setting $a(x)=a_{0}(x)+a_{1}(x)$, and likewise for $b(x)$, $s(x)$ and $e(x)$ we have that the even-degree part of $b(x)$ is
@@
  • set $count:=0$
  • do
    • $(a(x),b(x))\overset{X}{\leftarrow}R_{q}^{2}$
    • $count\leftarrow count+1$
  • until $a(x)\in R_{q,0}$
  • return $(a(x),b(x)),count$
@@
Assume now that $X$ is a PLWE distribution. Any $v=(a(x),b(x))$ sampled from $X$ has $a(x)$ sampled from an $R_{q}$-valued uniform distribution and $b(x)=a(x)s(x)+e(x)$, where $e(x)$ is sampled from $\chi_{R_{q}}$, an $R_{q}$-valued discrete truncated Gaussian variable. Taking into account that for any $i\geq 1$, it holds that
$P[C_{i,v}]=P[X=v]=P[U=a(x)\cap\chi_{R_{q}}=e(x)],$
@@
Within $X_{0}$, the variable countcount stores the expected number of times one has to run $X$ before succeeding. Since this random variable is distributed as a geometric random variable of first kind with success probability $q^{-1}$, the expected number of times is just its mean $E[count]=q=\mathcal{O}(N^{2})$ (see for instance [9, Appendix A]).
@@
Next, we must choose $k$ so that $1-q\left(|\Sigma|/q\right)^{k}\geq\theta$, where $\theta$ is our desired success probability threshold.
@@
  • set $C:=\emptyset$, $count:=0$, $c:=0$
  • do
[…]
    • $C\leftarrow C\cup\{(a(x),b(x))\}$
    • $count\leftarrow count+c$
  • until $|C|=k$ or $count>l$
  • if $count\leq l$ then
    • set $\omega:=$ Algorithm.2(C)(C)
    • return $\omega$
  • else
    • return FAIL